OpenStudy (anonymous):

Prove the identity using half angle formula: 2cos^2 θ/2 secθ = secθ + 1

OpenStudy (anonymous):

think\[\sec \theta \] as \[1/\cos \theta \], and try to simplify the equation first, (make one side equal to zero)

OpenStudy (anonymous):

I was working on the left hand side

OpenStudy (anonymous):

the idea was to use the half angle formula. at some point I ended up with (1 + cos) secθ = secθ + 1 and then I got stuck

OpenStudy (anonymous):

divide both sides by secθ, and you will get 1+cos on left, and 1/secθ which is equal to 1/1/cos = cos

OpenStudy (anonymous):

**on the right

OpenStudy (anonymous):

thank you